OpenStudy (anonymous):

Identify which table shows a direct variation. Two tables. For each table the column headings are x and y. Table 1 has column one values of 1, 2, 3 and column 2 values of 19, 38, 57. Table 2 has column 1 values 5, 6, 7 and column 2 values 48, 72, 84. A. Table 1 only B. Table 2 only C. both tables D. neither table

OpenStudy (anonymous):

@Whitemonsterbunny17

OpenStudy (anonymous):

@Abhisar @jim_thompson5910

OpenStudy (anonymous):

ok i wont forget to pick best response this time

OpenStudy (anonymous):

hi danny your the one who made hero delete my post about fans

OpenStudy (anonymous):

@dannyrod2000

OpenStudy (anonymous):

hi danny your the one who made hero delete my post about fans

OpenStudy (anonymous):

wow dude really y?

OpenStudy (anonymous):

just y?

OpenStudy (anonymous):

@Whitemonsterbunny17 can you help me?

whitemonsterbunny17 (whitemonsterbunny17):

\(\huge\frak\color{magenta}{Be~Nice.}\)

OpenStudy (anonymous):

me or danny

whitemonsterbunny17 (whitemonsterbunny17):

And no, I suck at math.

whitemonsterbunny17 (whitemonsterbunny17):

Both.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

dannyrod2000 stfu lol jk

OpenStudy (dannyrod2000):

@Hero

OpenStudy (anonymous):

hi hai hey heyyy sup wasup

OpenStudy (anonymous):

really danny stop gosh mannnnn jeez its 11:33 where i am jeez dude.

OpenStudy (dannyrod2000):

9:04 :P

whitemonsterbunny17 (whitemonsterbunny17):

Please keep replies relevant to the question.

OpenStudy (anonymous):

im not pacific lol

OpenStudy (anonymous):

ok sry bunny

OpenStudy (anonymous):

Im in connecticut its 3 hours later than pacific

OpenStudy (anonymous):

lol kiddo

OpenStudy (anonymous):

ok does anyone know my question?

OpenStudy (anonymous):

conerr change your pic now jeez its gross wow

whitemonsterbunny17 (whitemonsterbunny17):

Tagging him a bunch of times is unnecessary. He'll come when he has time.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok does anyone know my question?

OpenStudy (anonymous):

i need help plz?

whitemonsterbunny17 (whitemonsterbunny17):

Maybe @Compassionate or @hartnn can help.

OpenStudy (anonymous):

@shyguybros @jim_thompson5910

OpenStudy (anonymous):

why would you help danny bunny y jeez

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

can anyone help me i need help in math plz anyone? @shyguybros @jim_thompson5910 @Melodysim

OpenStudy (anonymous):

can anyone help me plz?

whitemonsterbunny17 (whitemonsterbunny17):

Err... Let's see.. @ShadowLegendX

hero (hero):

A direct variation is represented by y = kx. What it means is that for every value x in a set, there exists a value, y, such that y/x = k.

hero (hero):

For every (x,y) pair, if when you divide y/x and arrive at the same value k, then you have a direct variation.

hero (hero):

So for the first table divide: 19/1 = ? 38/2 = ? 57/3 = ?

hero (hero):

If you get the same value for each division, then you have a direct variation.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

What is the value of y when x = 20? Equation y = 5.2x. Over top of a table with column 1 heading x and values 0, 1, 10, 20 and column 2 heading y and no values showing. A. 10.4 B. 25.2 C. 100.4 D. 104

OpenStudy (anonymous):

anyone?

whitemonsterbunny17 (whitemonsterbunny17):

\(\Large\cal\color{blueviolet}{Thank~you,}\) @Hero \(\Large\frak\color{purple}{You're~}\) \(\huge\frak\color{red}{^A}\color{gold}{w}\color{lime}{_e}\color{cyan}{s}\color{blue}{^o}\color{purple}{m}\color{blueviolet}{_e}\color{magenta}{!}\)

OpenStudy (anonymous):

he is!!!!!

hero (hero):

y = 5.2x is a direct variation equation of the form y = kx and k = 5.2 For this particular case, x = 20 so replace x with 20 to get y = 5.2(20) In other words, to find y simply multiply 5.2 by 20.

OpenStudy (anonymous):

wow your good

hero (hero):

@lankymonkey you're welcome.

OpenStudy (anonymous):

your really smart what grade are you in

hero (hero):

I'm in the Infinite Grade

whitemonsterbunny17 (whitemonsterbunny17):

^ Infinitely Awesome Grade ∞